OpenStudy (anonymous):

I need help! I need to solve the integral of cos(ax)sin(bx), but I can't get it to what the answer is. Help! :(

OpenStudy (jamesj):

Let \[ I = \int \cos(ax) \sin(bx) \ dx \] and integrate twice by parts until you get the integral I back. In other words, you will obtain an expression that looks like this \[ I = (expression \ involving \ cos \ and \ sin ) + (some \ constant) I \] Then solve algebraically for \(I\).

OpenStudy (anonymous):

At the end though, wouldn't it be I = (expression) - (some constant)I? I don't want to get confused...haha That's the only difference I have in my work for it

OpenStudy (jamesj):

C'mon. That's a trivial difference, as the question is whether or not the constant has the -1 or not. +C = -(-C)

OpenStudy (anonymous):

Okay, okay. Just asking.

OpenStudy (jamesj):

So integrating once \[ I = 1/a . \sin ax . \sin bx - b/a \int \sin ax . \cos bx \ dx \] Now integrate again.

OpenStudy (anonymous):

Okay, I have that done. It just gets messy after that